# Max–Min Fairness in Stacked Intelligent Metasurface-Aided Rate Splitting Networks Abdullah Quran, Shimaa Naser, Maryam Tariq, Omar Alhussein, and Sami Muhaidat

### v1

@@ Abstract
This paper investigates a downlink multiuser multiple-input single-output system that integrates rate-splitting multiple access (RSMA) with a stacked intelligent metasurface (SIM) to enable wave-domain beamforming. Unlike conventional digital beamforming, the proposed system leverages the programmable phase shifts of the SIM to perform beamforming entirely in the wave domain. In contrast to existing literature, this work introduces a fairness-centric SIM-RSMA design that shifts the emphasis from maximizing sum-rate to ensuring fair allocation of resources. In particular, we formulate a max–min rate optimization problem that jointly optimizes transmit power coefficients at the base station and SIM phase shifts. Given the non-convex nature of this problem, we develop an alternating optimization framework, where the power allocation is optimized through successive convex approximation and SIM beamforming is optimized using the Riemannian conjugate gradient method. Simulation results indicate that combining SIM with RSMA yields superior max–min performance compared to its integration with space division multiple access or non-orthogonal multiple access.
Index Terms:
Rate-splitting multiple access (RSMA), stacked intelligent metasurface (SIM), wave-domain beamforming.
@@ I Introduction
With the increasing demand for higher data rates and more energy-efficient infrastructure, traditional beamforming techniques that rely heavily on digital signal processing and radio frequency (RF) chains become increasingly complex, and costly, particularly for multi-user environments. In this regard, reconfigurable intelligent surfaces (RISs) have emerged as a transformative technology, enabling programmable control over electromagnetic (EM) wave propagation via passive meta-atom arrays [1]. These surfaces enhance spectral efficiency and energy efficiency by imposing tunable phase shifts on incident waves [2, 3]. However, conventional RISs are typically limited to single-layer reflection, acting as passive relays with limited spatial processing capabilities, which constrains their performance in complex wireless environments. To address these challenges, stacked intelligent metasurfaces (SIMs), inspired by multi-layer diffractive optics [4], have been recently proposed to enable complex beamforming operations in the wave domains using multiple stacked metasurface layers, which reduces the reliance on digital beamforming hardware [5, 6]. Recent works have demonstrated the potential of SIMs in multiuser wave-domain beamforming [7, 8], and integrated sensing and communication [9].
@@
Likewise, rate-splitting multiple access (RSMA) has emerged as a powerful and flexible strategy for multiuser interference management. By splitting each message into a common part decoded by all users and private parts decoded individually, RSMA enables partial interference decoding yielding superior spectral efficiency compared to conventional schemes such as space-division multiple access (SDMA) and non-orthogonal multiple access (NOMA) [10, 11, 12].
@@
Recent studies have explored the integration of RSMA with RISs to enhance spectral efficiency and interference management [13]. However, their performance is constrained by the limited beamforming resolution of single-layer RIS. In contrast, SIM enables high-precision beamforming, suppressing inter-user interference via focused beam alignment. This fine-grained spatial control complements the interference management strategy of RSMA, which mitigates residual interference through message splitting. When combined, SIM and RSMA form a robust and synergistic framework for improving multiuser performance in interference-limited and dense wireless environments. Furthermore, SIM’s analog architecture reduces energy consumption and hardware complexity by eliminating the need for a digital beamformer. Despite these advantages, the integration of SIM and RSMA remains unexplored, particularly in fairness-centric designs such as max–min rate optimization.
Existing SIM literature primarily focuses on single-user systems [5] or conventional multiple access schemes [8]. While RSMA has been considered in metasurface-assisted systems where dual-polarized SIMs are used to enhance digital beamforming [14], this hybrid analog–digital architecture differs from the fully wave-domain beamforming approach adopted in this work. At the time of submission, the authors became aware of an early-access paper [15], which is the only existing work that explores SIM-RSMA systems and focuses on sum-rate maximization. However, it does not address fairness, a major challenge in multiuser systems where a desired quality of service must be ensured for all users. To bridge this gap, we propose a fairness-centric SIM-assisted RSMA framework for downlink multiuser MISO systems, wherein SIM performs all beamforming operations entirely in the wave domain. This design eliminates the need for digital precoding hardware while enabling efficient and equitable user performance. The main contributions of this work are summarized as follows:
  • We propose a SIM-RSMA scheme where wave-domain beamforming via the SIM synergizes with RSMA’s message splitting to maximize the minimum user rate. To this end, a max–min rate optimization problem is formulated to jointly optimize the transmit power and SIM phase shifts. However, solving the formulated problem is inherently difficult due to strong coupling between power allocation variables and SIM phase shifts within the non-convex objective function and constraints, making conventional optimization approaches inadequate.
@@
  • To address the non-convexity arising from the coupling between power allocation and SIM phase shifts, an alternating optimization (AO) framework is proposed. Transmit power is updated using successive convex approximation (SCA), while SIM beamforming is refined using the Riemannian conjugate gradient (RCG) method.
@@
  • The proposed SIM-RSMA scheme performance is evaluated against SIM-SDMA and SIM-NOMA baselines. Results show that SIM-RSMA achieves a higher minimum user rate, demonstrating the effectiveness of combining wavedomain beamforming with RSMA’s interference management. Moreover, the use of multilayer SIM is shown to achieve better fairness performance compared to conventional single-layer RIS structures
@@ II System Model and Problem Formulation
As illustrated in Fig. 1, we consider a downlink multiuser MISO system where a base station (BS) equipped with $N$ antennas communicates with $K$ single-antenna user equipments (UEs) via 1-layer RSMA [11]. In this work, we propose a SIM-based architecture. The SIM consists of $L$ programmable metasurface layers with $M$ meta-atoms per layer, to perform beamforming directly in the EM domain. The transmitted signal consists of a common stream decodable by all UEs and private streams intended for individual UEs. Let =$\mathcal{L}\!\!=\!\!\{1,2,\ldots,L\}$, =$\mathcal{M}\!\!=\!\!\{1,2,\ldots,M\}$, and $\mathcal{K}\!\!=\!\!\{1,2,\ldots,K\}$ represent the sets of metasurfaces, meta-atoms per layer, and UEs, respectively. The phase shift matrix corresponding to the $\ell$-th metasurface layer is denoted as $\mathbf{\Theta}_{\ell}\!\!=\!\!\text{diag}\left(\theta_{\ell,1},\theta_{\ell,2},\ldots,\theta_{\ell,M}\right)\in\mathbb{C}^{M\times M}$, $\forall\ell\!\!\in\!\!\mathcal{L}$, where $\theta_{\ell,m}\!\!=\!\!e^{j\phi_{m}^{\ell}}$ represents the EM response of the $m$-th meta-atom on the $\ell$-th metasurface layer, and $\phi_{m}^{\ell}\in[0,2\pi)$ denotes the corresponding phase shift.
@@
The Rayleigh–Sommerfeld diffraction theory is employed to characterize the EM coupling between adjacent metasurface layers. Specifically, the propagation between the $(\ell-1)$-th and $\ell$-th metasurface layers is described by a transmission matrix $\mathbf{Q}_{\ell}\in\mathbb{C}^{M\times M}$, defined for all $\ell\in\mathcal{L}$, $\ell\neq 1$. The $(m,m^{\prime})$-th entry, $q_{\ell}^{m,m^{\prime}}$, of $\mathbf{Q}_{\ell}$ denotes the transmission coefficient from the $m^{\prime}$-th meta-atom on the $(\ell-1)$-th layer to the $m$-th meta-atom on the $\ell$-th layer, and is given by
$\footnotesize q_{\ell}^{m,m^{\prime}}=\frac{A_{t}\cos\psi_{\ell}^{m,m^{\prime}}}{d_{\ell}^{m,m^{\prime}}}\left(\frac{1}{2\pi d_{\ell}^{m,m^{\prime}}}-j\frac{1}{\lambda}\right)e^{j2\pi d_{\ell}^{m,m^{\prime}}/\lambda},$ (1)
@@
where $\lambda$ is the wavelength, $A_{t}$ is the area of each meta-atom, $\psi_{\ell}^{m,m^{\prime}}$ denotes the angle between the propagation direction and the surface normal, and $d_{\ell}^{m,m^{\prime}}$ is the corresponding propagation distance. The BS-to-SIM transformation matrix is denoted by $\mathbf{Q}_{1}=[\mathbf{q}_{1},\mathbf{q}_{2},\ldots,\mathbf{q}_{N}]\in\mathbb{C}^{M\times N}$, where each $\mathbf{q}_{n}\in\mathbb{C}^{M\times 1}$ models the response from the $n$-th transmit antenna to the first metasurface layer. Following the convention in [9, 7, 8, 15], we assume each antenna is associated with one user. To support RSMA with $K$ private streams and one common stream, we set $N=K+1$, where the first $K$ antennas correspond to the private messages and the $(K+1)$-th antenna to the common message. The complete wave-domain response of the SIM is expressed as
$\footnotesize\mathbf{F}=\mathbf{\Theta}_{L}\mathbf{Q}_{L}\mathbf{\Theta}_{L-1}\mathbf{Q}_{L-1}\cdots\mathbf{\Theta}_{2}\mathbf{Q}_{2}\mathbf{\Theta}_{1}\in\mathbb{C}^{M\times M}.$ (2)
@@
To model the end-to-end downlink channel from the SIM to each user, we assume a quasi-static flat-fading scenario. Let $\mathbf{h}_{k}\in\mathbb{C}^{M\times 1}$ denote the effective channel vector from the SIM to the $k$-th user. The channel $\mathbf{h}_{k}$ is assumed to follow a correlated Rayleigh fading model given by $\mathbf{h}_{k}\sim\mathcal{CN}(\mathbf{0},\zeta_{k}\mathbf{T})$, where $\zeta_{k}$ represents the distance-dependent path loss, and $\mathbf{T}\in\mathbb{C}^{M\times M}$ is the spatial correlation matrix characterizing coupling between the meta-atoms. Under an isotropic multipath environment with a uniform power azimuth spectrum, the correlation between the $n$-th and $n^{\prime}$-th meta-atoms is modeled as $[\mathbf{T}]_{n,n^{\prime}}=\text{sinc}\left(\frac{2d_{n,n^{\prime}}}{\lambda}\right)$, where $d_{n,n^{\prime}}$ is the spacing between meta-atoms, $\lambda$ is the carrier wavelength, and $\text{sinc}(x)=\sin(\pi x)/(\pi x)$ is the normalized sinc function [16].
@@
Using the 1-layer RSMA strategy [11], the BS transmits a signal that can be expressed as $\mathbf{x}=\sqrt{p_{c}}s_{c}+\sum_{k\in\mathcal{K}}\sqrt{p_{k}}s_{k},$ where $s_{c}\sim\mathcal{CN}(0,1)$ is the common message decoded by all users, and $s_{k}\sim\mathcal{CN}(0,1)$ is the private message intended for user $k$. Accordingly, the received signal at user $k$ is given by
$\footnotesize y_{k}=\mathbf{h}_{k}^{H}\mathbf{F}\left(\mathbf{q}_{c}\sqrt{p_{c}}s_{c}+\sum_{j\in\mathcal{K}}\mathbf{q}_{j}\sqrt{p_{j}}s_{j}\right)+n_{k},$ (3)
@@
At the receiving side, each user applies successive interference cancellation (SIC) to decode its messages. The signal-to-interference-plus-noise ratios (SINRs) for decoding the common and private messages at user $k$ are given by
$\footnotesize\gamma_{\delta,k}=\frac{p_{\delta}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{\delta}|^{2}}{\sum_{j\in\mathcal{K}_{\delta}(k)}p_{j}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}+\sigma_{k}^{2}},\quad\delta\in\{\mathrm{c},\mathrm{p}\}$ (4)
where $\delta\in\{\mathrm{c},\mathrm{p}\}$, and the index sets are $\mathcal{K}_{\mathrm{c}}(k)=\mathcal{K}$ and $\mathcal{K}_{\mathrm{p}}(k)=\mathcal{K}\setminus\{k\}$. Accordingly, the achievable common and private rates for user $k$ are given by $R_{c,k}=\log_{2}(1+\gamma_{c,k})$ and $R_{p,k}=\log_{2}(1+\gamma_{p,k})$, respectively. Since all users must successfully decode the common message, its rate is constrained by the weakest user, i.e., $R_{c}\leq\min_{k}R_{c,k}$. Consequently, the individual achievable rate is given by $R_{k}=R_{c,k}+R_{p,k}$.
Aiming to maximize the minimum rate across all users, we formulate the joint optimization of the transmit power allocation and the phase shift configuration of the SIM as follows:
@@
$\quad p_{c}+\sum_{k\in\mathcal{K}}p_{k}\leq P_{\max},$ (5c)
$\quad\sum_{k\in\mathcal{K}}r_{c,k}\leq R_{c,k},\quad r_{c,k}\geq 0,\quad\forall k\in\mathcal{K}$ (5d)
where $\mathbf{p}=[p_{c},p_{1},\ldots,p_{K}]$ denotes the vector of power allocation variables and $\mathbf{r}_{c}=[r_{c,1},r_{c,2},\ldots,r_{c,K}]$ denotes the vector of auxiliary variables representing the common rate portions assigned to the users. Furthermore, constraint (5b) imposes a unit-modulus condition on each phase shift element of the SIM. On the other hand, constraint (5c) limits the total transmit power at the BS. Finally, constraint (5d) guarantees that the sum of the allocated common rates across all users remains within the decodable common rate at each user. Problem (P1) is a non-convex and challenging optimization problem due to the coupled design of the transmit power variables and the SIM phase shifts in (5a), as well as the non-convex unit-modulus constraints in (5b). Therefore, obtaining the globally optimal solution to (P1) is intractable. To address this, we propose an efficient algorithm based on AO to obtain a near-optimal solution for (P1).

### v2

@@ Abstract
This paper investigates a downlink multiuser multiple-input single-output system that integrates rate-splitting multiple access (RSMA) with a stacked intelligent metasurface (SIM) to enable wave-domain beamforming. Unlike conventional digital beamforming, the proposed system leverages the programmable phase shifts of the SIM to perform beamforming entirely in the wave domain. This work introduces a fairness-centric SIM-RSMA design that ensures equitable resource allocation, which is critical for applications such as industrial IoT and multi-user extended reality. We formulate a max-min rate optimization problem that jointly optimizes the base station’s transmit power and the SIM’s phase shifts. Due to the non-convex and coupled nature of this problem, we develop an alternating optimization framework where the power allocation is optimized via successive convex approximation and the SIM beamforming is optimized using the Riemannian conjugate gradient method. Simulation results demonstrate that the proposed SIM-RSMA framework achieves superior minimum user rates and enhanced fairness compared to SIM-assisted non-orthogonal multiple access schemes.
@@ I Introduction
With the increasing demand for higher data rates and more energy-efficient communication infrastructure, traditional beamforming techniques rely heavily on digital signal processing and radio frequency (RF) chains are complex and costly for multiuser environments. Reconfigurable intelligent surfaces (RISs) have emerged as a transformative technology, enabling programmable control over electromagnetic (EM) wave propagation via passive meta-atom arrays [ris_lit1]. RISs enhance spectral efficiency and energy efficiency by applying tunable phase shifts on incident waves [ris_lit6]. However, conventional RISs are typically limited to single-layer reflection, acting as passive relays with limited spatial processing capabilities, which constrains their performance in complex wireless environments. To address these challenges, stacked intelligent metasurfaces (SIMs), inspired by multi-layer diffractive optics [sim_lit2], have been recently proposed to enable complex beamforming operations in the wave domain using multiple stacked metasurface layers, thereby reducing the reliance on digital beamforming hardware [sim_lit3, sim_lit4].
@@
Likewise, rate-splitting multiple access (RSMA [rsma0]) has emerged as a powerful and flexible strategy for multiuser interference management. RSMA splits each message into a common part decoded by all users and private parts decoded individual users. Hence, RSMA enables partial interference decoding, yielding superior spectral efficiency compared to conventional schemes such as space-division multiple access (SDMA [sdma]) and non-orthogonal multiple access (NOMA [rsma0])
@@
The integration of RSMA with SIM presents a promising direction for next-generation wireless networks. SIM enables high-precision beamforming, suppressing inter-user interference via focused beam alignment. This fine-grained spatial control complements the interference management strategy of RSMA, which mitigates residual interference through message splitting. Furthermore, SIM’s analog architecture reduces energy consumption and hardware complexity by eliminating the need for digital beamforming. Combining SIM and RSMA form a robust and synergistic framework for improving performance and reducing cost and power in multiuser interference-limited and dense wireless environments.
Despite these advantages, the integration of SIM and RSMA remains unexplored. Recent studies have either explored the integration of RSMA with RISs [IRS_RSMA1], which is constrained by the limited beamforming resolution of single-layer RISs, or focused on SIM for single-user systems and conventional multiple access [sim_lit3, sim_multi_user2]. A notable exception is the work presented in [sim_rsma] which considers SIM-RSMA in the wave domain but focuses on sum-rate maximization. This approach has critical limitations for multiuser systems where ensuring a uniform quality of service (QoS) is paramount, such as in industrial IoT sensor networks, autonomous vehicles, and multiuser extended reality applications [fairness_applications]. In these scenarios, the sum-rate objective fails to ensure rate allocation, often collapsing to a trivial solution that allocates most resources to users with the strongest channels, leading to resource deprivation for the remaining users.
In this work, we propose a max-min fairness framework for wave-domain SIM-RSMA systems that ensures balanced performance across all users. The proposed design effectively combines SIM’s fine-grained beamforming with RSMA’s interference management to enhance robustness and guarantee uniform rate QoS in multiuser environments. However, achieving such fairness is significantly more challenging than sum-rate oriented designs as the optimization problem becomes highly non-convex and tightly coupled, involving joint power allocation and multi-layer phase-shift beamforming under strict unit-modulus constraints. Furthermore, the beamforming design becomes more sophisticated, requiring the SIM to prioritize users with weak channels and manage interference inclusively. This leads to interference-rich conditions that fully exploit RSMA’s potential: the common stream mitigates interference for all users, while the SIM’s beamforming focuses private streams to reduce mutual interference. The main contributions of this work are summarized as follows:
  • We propose the first max-min fairness design for the SIM-RSMA system, where wave-domain beamforming and RSMA-based message splitting are jointly optimized to maximize the minimum user rate. The formulated problem is highly non-convex due to coupled power and phase-shift variables and unit-modulus constraints.
@@
  • We develop an efficient alternating optimization (AO) framework to address this challenge. The power allocation and common rate are optimized via successive convex approximation (SCA), while the SIM phase shifts are updated using Riemannian conjugate gradient (RCG) on the complex circle manifold.
@@
  • We evaluate the proposed SIM-RSMA scheme against SIM-SDMA and SIM-NOMA baselines. Results show that SIM-RSMA outperforms SIM-NOMA while achieving comparable performance to SIM-SDMA in terms of minimum user rate and fairness. Moreover, the use of multilayer SIM is shown to achieve better performance compared to conventional single-layer SIM structures.
@@ II System Model and Problem Formulation
As illustrated in Fig. 1, we consider a downlink multiuser MISO system where a base station (BS) equipped with $N$ antennas communicates with $K$ single-antenna user equipments (UEs) via RSMA. In this work, we propose a SIM-based architecture to perform beamforming directly in the wave domain. The SIM consists of $L$ programmable metasurface layers with $M$ meta-atoms per layer to perform beamforming directly in the EM domain. The transmitted signal consists of a common stream decodable by all UEs and private streams intended for individual UEs. Let $\mathcal{L}\!\!=\!\!\{1,2,\ldots,L\}$, $\mathcal{M}\!\!=\!\!\{1,2,\ldots,M\}$, and $\mathcal{K}\!\!=\!\!\{1,2,\ldots,K\}$ represent the sets of metasurfaces, meta-atoms per layer, and UEs, respectively. The phase shift matrix corresponding to the $\ell$-th metasurface layer is denoted as $\mathbf{\Theta}_{\ell}\!\!=\!\!\text{diag}\left(\theta_{\ell,1},\theta_{\ell,2},\ldots,\theta_{\ell,M}\right)\in\mathbb{C}^{M\times M}$, $\forall\ell\!\!\in\!\!\mathcal{L}$, where $\theta_{\ell,m}\!\!=\!\!e^{j\phi_{m}^{\ell}}$ represents the EM response of the $m$-th meta-atom on the $\ell$-th metasurface layer, and $\phi_{m}^{\ell}\in[0,2\pi)$ denotes the corresponding phase shift.
@@
The Rayleigh–Sommerfeld diffraction model [sim_lit2] is employed to capture the near-field wave interactions between the metasurface layers, which is essential for modeling the wave-based beamforming capabilities of the multi-layer SIM. Specifically, the wave propagation from the $(\ell-1)$-th to the $\ell$-th metasurface layer is modeled by a transmission matrix $\mathbf{Q}_{\ell}\in\mathbb{C}^{M\times M}$, defined for all $\ell\in\mathcal{L}$, $\ell\neq 1$. The $(m,m^{\prime})$-th entry, $q_{\ell}^{m,m^{\prime}}$, of $\mathbf{Q}_{\ell}$ represents the transmission coefficient from the $m^{\prime}$-th meta-atom on the $(\ell-1)$-th layer to the $m$-th meta-atom on the $\ell$-th layer, and is given by
$q_{\ell}^{m,m^{\prime}}=\frac{A_{t}\cos\psi_{\ell}^{m,m^{\prime}}}{d_{\ell}^{m,m^{\prime}}}\left(\frac{1}{2\pi d_{\ell}^{m,m^{\prime}}}-j\frac{1}{\lambda}\right)e^{j2\pi d_{\ell}^{m,m^{\prime}}/\lambda},$ (1)
@@
where $\lambda$ is the wavelength, $A_{t}$ is the area of each meta-atom, $\psi_{\ell}^{m,m^{\prime}}$ is the angle between the propagation direction and the surface normal, and $d_{\ell}^{m,m^{\prime}}$ is the corresponding propagation distance. The BS-to-SIM transformation matrix is denoted by $\mathbf{Q}_{1}=[\mathbf{q}_{1},\mathbf{q}_{2},\ldots,\mathbf{q}_{N}]\in\mathbb{C}^{M\times N}$, where each $\mathbf{q}_{n}\in\mathbb{C}^{M\times 1}$ models the response from the $n$-th transmit antenna to the first metasurface layer. Following the convention in [sim_isac, sim_multi_user, sim_multi_user2, sim_sum_rate], we assume each antenna is associated with one user. To implement 1-layer RSMA, the BS is equipped with $N=K+1$ antennas: $K$ antennas transmit the private streams and the remaining antenna transmit the common stream. This configuration ensures independent control over both the transmit power and the achievable signal-to-interference-plus-noise ratio (SINR) for all stream, while granting the SIM sufficient spatial flexibility to manage multi-user interference through wave-based beamforming. The complete wave-domain response of the SIM is expressed as
$\mathbf{F}=\mathbf{\Theta}_{L}\mathbf{Q}_{L}\mathbf{\Theta}_{L-1}\mathbf{Q}_{L-1}\cdots\mathbf{\Theta}_{2}\mathbf{Q}_{2}\mathbf{\Theta}_{1}\in\mathbb{C}^{M\times M}.$ (2)
@@
To model the end-to-end downlink channel from the SIM to each user, we assume a quasi-static flat-fading scenario. Let $\mathbf{h}_{k}\in\mathbb{C}^{M\times 1}$ denote the effective channel vector from the SIM to the $k$-th user. The channel $\mathbf{h}_{k}$ is assumed to follow a correlated Rayleigh fading model given by $\mathbf{h}_{k}\sim\mathcal{CN}(\mathbf{0},\zeta_{k}\mathbf{T})$, where $\zeta_{k}$ represents the distance-dependent path loss and $\mathbf{T}\in\mathbb{C}^{M\times M}$ is the spatial correlation matrix characterizing coupling between the meta-atoms. Under an isotropic multipath environment with a uniform power azimuth spectrum, the correlation between meta-atoms is represented by the Fourier transform of the uniform angular distribution. Thus, the correlation between the $m$-th and $m^{\prime}$-th meta-atoms is modeled as $\mathbf{T}]_{m,m^{\prime}}=\text{sinc}\left(2d_{m,m^{\prime}}/\lambda\right)$, where $d_{m,m^{\prime}}$ is the spacing between meta-atoms [correlation]. Moreover, we assume perfect channel state information (CSI) of the SIM-User channels $\mathbf{h}_{k}$ is available at the BS.
@@
Using the 1-layer RSMA strategy [rsma0], the BS transmits a signal that can be expressed as $\mathbf{x}=\sqrt{p_{c}}s_{c}+\sum_{k\in\mathcal{K}}\sqrt{p_{k}}s_{k},$ where $s_{c}\sim\mathcal{CN}(0,1)$ is the common message decoded by all users, and $s_{k}\sim\mathcal{CN}(0,1)$ is the private message intended for user $k$. Accordingly, the received signal at user $k$ is given by
$y_{k}=\mathbf{h}_{k}^{H}\mathbf{F}\left(\mathbf{q}_{c}\sqrt{p_{c}}s_{c}+\sum_{j\in\mathcal{K}}\mathbf{q}_{j}\sqrt{p_{j}}s_{j}\right)+n_{k},$ (3)
@@
At the receiving side, each user applies successive interference cancellation (SIC) to decode its messages. The SINRs for decoding the common and private messages at user $k$ are given by
$\gamma_{\delta,k}=\frac{p_{\delta}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{\delta}|^{2}}{\sum_{j\in\mathcal{K}_{\delta}(k)}p_{j}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}+\sigma_{k}^{2}},\quad\delta\in\{\mathrm{c},\mathrm{p}\}$ (4)
where $\delta\in\{\mathrm{c},\mathrm{p}\}$, and the index sets are $\mathcal{K}_{\mathrm{c}}(k)=\mathcal{K}$ and $\mathcal{K}_{\mathrm{p}}(k)=\mathcal{K}\setminus\{k\}$. Accordingly, the achievable common and private rates for user $k$ are given by $R_{c,k}=\log_{2}(1+\gamma_{c,k})$ and $R_{p,k}=\log_{2}(1+\gamma_{p,k})$, respectively. Since all users must successfully decode the common message, its rate is constrained by the weakest user, i.e., $R_{c}\leq\min_{k}R_{c,k}$. Consequently, the individual achievable rate is given by $R_{k}=r_{c,k}+R_{p,k}$ such that $\sum_{k\in\mathcal{K}}r_{c,k}\leq R_{c}$, where $\mathbf{r}_{c}=[r_{c,1},r_{c,2},\ldots,r_{c,K}]$ denotes the vector of auxiliary variables representing the common rate portions assigned to the users.
Aiming to maximize the minimum rate across all users, we formulate the joint optimization of the transmit power allocation and the phase shift configuration of the SIM as
$(\text{P1}):$ $\max_{\mathbf{\Theta}_{l},\mathbf{p},\mathbf{r}_{c}}\quad\min_{k\in\mathcal{K}}\{r_{c,k}+R_{p,k}\}$ (5a)
$\quad\text{s.t.}\quad|\theta_{\ell,m}|=1,\quad\quad\quad\quad\forall\ell\in\mathcal{L},\forall m\in\mathcal{M},$ (5b)
@@
where $\mathbf{p}=[p_{c},p_{1},\ldots,p_{K}]$ is the vector of power allocation variables. Constraint (5b) imposes a unit-modulus condition on each phase shift element of the SIM. constraint (5c) imposes a constraint on the power transmitted by the BS, ensuring it does not exceed $P_{\max}$, which represents the maximum power budget at the BS. Finally, constraint (5d) guarantees that the sum of the allocated common rates across all users remains within the decodable common rate at each user.